What is the answer and steps to solving this problem
1/15(2x+5)=x+2/9

Answer by stanbon(75887)

1/15(2x+5)= (x+2)/9
---------
Multiply both sides by 9*15 = 135
-----
9(2x+5) = 15(x+2)
----------------
18x + 45 = 15x + 30
3x = -15
x = -5
-----------------------
Cheers,
Stan H.
=======================